January 3, 1990 Anderson Company sells a house with a cost of $700,000,000 for sale at $1,000,000,000 at 10% interest per year. installments are made every semester for 5 years (10x installments) and 20% down payment calculated from the declining balance. Journalize all transactions in which gross profit is recognized in proportion to the collection per cash.


1
Expert's answer
2022-01-10T09:53:56-0500

Debit of accounts Cash Register, Settlement account, Settlements with customers credit Sales account - 200 000 000

1 000 000 000-200 000 000=800 000 000

This is an annuity

i=0.05

n=10

"A=800 000 000\frac{0.05}{1-(1+0.05)^{-10}}=103 603 659.97"

Debit of accounts Cash Register, Settlement account, Settlements with customers credit Sales account - 103 603 659.97